how do I solve this problem for radical equations.
1. (x-2)to the 1/3 power=5 Check for extraneous solution.

Answer by stanbon(48558)

(x-2)to the 1/3 power=5
-----------
(x-2)^(1/3) = 5
Cube both sides to get:
x-2 = 125
x = 127
============
Cheers,
Stan H.